OpenStudy (anonymous):

WILL FANN & MEDALL How many real-number solutions does the equation have? -7x^2 + 6x + 3 = 0 a. one solution b. two solutions c. no solutions d. infinitely many solutions

OpenStudy (anonymous):

@Here_to_Help15

OpenStudy (anonymous):

Let first check the discriminant value of this equation shall we: \[ax^2 + bx + c = 0\] Discriminant formula says that: \[\Delta = b^2 + 4ac\] if \[\Delta = 0\] there wil be one real-value solution. if \[\Delta < 0\] there will be no real-value solution if \[\Delta > 0\] There will be two real value solution. In here \[\Delta = 6^2 - 4 \times (-7)(3) = 120 > 0\] So as we stated earlier discriminant is greater than 0. So two real values solution exist.

OpenStudy (anonymous):

thxxx!!! :)